OpenStudy (anonymous):

isolate y for the following equation:

OpenStudy (anonymous):

\[x= -14e ^{2-y} + y^3 + 7y + 6\]

OpenStudy (anonymous):

i don't think there is an elementary method to do that.

OpenStudy (anonymous):

hmm the original equation asks to find the inverse function of \[-14e ^{2-x} + x^3 + 7x + 6 \]

OpenStudy (anonymous):

which means swapping y for x and solving for y

OpenStudy (anonymous):

and then find the equation of the line perpendicular to this inverse function at x = 14

OpenStudy (anonymous):

correct... but isolating the y after you reversed it is very difficult. i guess all you could do here is give the function implicitly.

OpenStudy (anonymous):

can you give the original problem? looks like you need to take derivative implicitly here.

OpenStudy (anonymous):

differentiate both sides of the equation, find y' maybe

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let \[f(x) = -14e ^{2-x} + x^3 + 7x + 6\] find an equation of the normal line to the graph of f-1(x) at the point on the graph where x=14

OpenStudy (anonymous):

normal line btw, means the line w/ a negative reciprocal

OpenStudy (anonymous):

aka perpendicular. And the answer to this is y = -33x+464

OpenStudy (anonymous):

do you know how the derivatives (slopes) of inverse functions are related?

OpenStudy (anonymous):

maybe could you state it?

OpenStudy (anonymous):

i get \[\frac{ dy }{ dx } = \frac{ 1 }{ 14e ^{2-y} + 3y^2 + 7 }\]

OpenStudy (anonymous):

my prof gave a hint saying we had to find a value for x that would make the \[e ^{2-x} \] part equal to 0. dont know if that helps tho

hartnn (hartnn):

no, need to find the inverse function, u need to find y when x=14 lets say y=2 so u get \(f^{-1}(14)=2\) then (14,2) is on the graph of x = f-1(y). The slope of that function at that point is what you are looking for.

OpenStudy (anonymous):

how do we find\[f ^{-1}(14)\]

OpenStudy (anonymous):

since f(2)=14, then f-1(14)=2

OpenStudy (anonymous):

oh right

OpenStudy (anonymous):

wasnt 2 just used as a hypothetical

hartnn (hartnn):

yeah, thats a challenge \(14 = -14e ^{2-x} + x^3 + 7x + 6\) solve for x to get x=2 is not easy

hartnn (hartnn):

i found it by trial and error method

OpenStudy (anonymous):

oh ok

hartnn (hartnn):

the 1st thing i thought is to eliminate e, by putting x=2 and voila!

OpenStudy (anonymous):

oh i see

hartnn (hartnn):

so now u have a point, the derivative,can u find the slope of line....and then the equation?

OpenStudy (anonymous):

yes that makes it way easier

OpenStudy (anonymous):

thanks to both of you

hartnn (hartnn):

*point and derivative of inverse function....

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

yw... also remember you're not looking for the tangent but the normal line

hartnn (hartnn):

i got this because i have seen a similar question here, http://www.physicsforums.com/showthread.php?t=397539

OpenStudy (anonymous):

i see